OpenStudy (anonymous):

need a check A DOG IS FIRED VERTICALLY INTO THE AIR FROM A CANNON, WITH A VELOCITY OF 146FT/SEC. what is the altitude of the dog after 4 sec. and after 5.125 secs.? I GOT 4sec=584ft and 5.125sec.=748 ft

OpenStudy (amistre64):

ft means -16t^2 initial velocity is 146; and initial height is ...negligible

OpenStudy (amistre64):

\[h(t) = -16t^2+146t\] right?

OpenStudy (anonymous):

my notes gives me the formula 16t^2+Vot+So thats the same thing rite? u just plugged 0 for So?

OpenStudy (amistre64):

I plugged 0 for So; but your notes need to have the "-16t^2" or else you get wrong results

OpenStudy (amistre64):

at t=4; this is 328

OpenStudy (anonymous):

ohhh i forgot the negative!!!

OpenStudy (amistre64):

yeah, it helps when gravity actually pulls you back ;)

OpenStudy (anonymous):

once again you've saved the day!!! :)

OpenStudy (amistre64):

the value I get for 5.125 is 328 also; but I dont know if that means we hit the bend, or if its to small for my calculator to go any further

OpenStudy (anonymous):

thats the same thing i got and im using a TI-84..DOES THAT MEAN MY ANSWER IF OFF??

OpenStudy (amistre64):

the answer you gave above is off :) but I just graphed it on me ti83plus and it shows its just on opposite sides of the bend; so its good 328 is the answer for both t=4 and t=5.125

OpenStudy (anonymous):

OK THANKS!!!

OpenStudy (amistre64):

one way to tell is to see if the vertex is exactly between 4 and 5.125 -32t +146 = 0 at t = 146/32 146/32 -n = 4 146/32 +n = 5.125 ------------------ 146/16 = 9.125 ? its true

OpenStudy (anonymous):

ITS ALSO ASK HOW LONG UNTIL THE DOG LANDS I DNT EVN KNOW HOW TO BEGIN THAT ONE?

OpenStudy (amistre64):

when the dog lands; h = 0 again;

OpenStudy (amistre64):

but, the vertex tells us how long half the trip takes :) so, landing time will take twice as long as it takes to reach the vertex. vertex is at t = 146/32 right? now double it

OpenStudy (amistre64):

should be 9.125 ...

OpenStudy (anonymous):

i dnt understand? u just added the two times?

OpenStudy (amistre64):

dbk chk that tho

OpenStudy (amistre64):

not really adding the 2 times but we know there is no forward momentum, since the dog is shot straight up; so the time it takes to reach the full height is the same amount of time it takes to fall back down to the ground right?

OpenStudy (amistre64):

the 'vertex' is the point where it reaches it full height; so just double the time it takes to reach the vertex

OpenStudy (amistre64):

since the vertex is half way between 4 and 5.125; add them together and divide by 2; then double it

OpenStudy (amistre64):

\[\frac{4+5.125}{2}*2 = 4+5.125=9.125\]

OpenStudy (anonymous):

ok thanx agin! well off to this algebra class!:( ttyl!!

OpenStudy (amistre64):

Ciao bella ;)

OpenStudy (anonymous):

Mon dieu, poor dog.